The three most prominent buildings in a city, Washington Center, Lincoln Galleria, and Jefferson Square Tower have a total height of 1800 feet. Find the height of each building if Jefferson Square Tower is twice as tall as Lincoln Galleria and Washington Center is 200feet taller than Lincoln Galleria.
Answer by Paul(988)

Let Lincoln be x
Let Square tower be 2x
Let Washington be 200+x
(x)+(2x)+(200+x)=1800
4x+200=1800
4x=1600
x=400

2(400)=800
200+400=600

Hence, the Lincoln is 400ft, the Square is 800ft, and the Washington is 600ft.
Paul.